OpenStudy (anonymous):

PLEASE HELP I WILL MEDAL! What is the standard form of equation of a circle with center at (-2, -3) and passing through the point (-2, 0)? A. (x + 2)^2 + (y + 3)^2 = 9 B. (x − 2)^2 + (y − 3)^2 = 16 C. (x − 2)^2 + (y − 3)^2 = 4 D. (x + 2)^2 + (y + 3)^2 = 16 E. (x − 2)^2 + (y + 3)^2 = 9

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

What's the distance between the points (-2, 0) and (-2, 3)? That distance will be the radius.

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

no, the x's are the same, so we can just subtract the y's. r = 0 - (-3) = 3

OpenStudy (anonymous):

now just plug into the formula (x - h)² + (y - k)² = r² (h, k) is (-2, -3) the center

OpenStudy (anonymous):

(x-2)^2 +(y-3)^2=9

OpenStudy (anonymous):

do i switch them to addition since they were already negatives?

OpenStudy (anonymous):

yes you have to switch them

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome